OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is -11, the last term is -45,056, and the common ratio is -4

OpenStudy (kirbykirby):

Geometric series: \[a+ar+ar^2+ar^3+... \] If there are 7 terms: \[\sum_{n=1}^7ar^{n-1}=a+ar+ar^2+...+ar^6 \] First term: \(a=-11\) Last term: \(ar^6=-45,056\) But \(r=-4\) So you can add these individually as: \(-11+(-11)(-4)+(-11)(-4)^2+...+(-11)(-4)^6\) Or the formula which is easier: \[ a\left(\frac{1-r^n}{1-r} \right)\]